OpenStudy (anonymous):

determine which equation below has a solution of x=2 a. 5x+8-12x=-6x+2x-3 b. 6x+4-12x=-10x+18x-3 c. -4x-6+x=-2x+x-10 d. -3x-8+x=-4x+x-10

OpenStudy (anonymous):

Try simplifying and making x the subject for all the equations, i.e.: 5x+8-12x=-6x+2x-3 5x-12x+6x-2x = -3-8 * bringing all the x to one side *

OpenStudy (anonymous):

You should first combine like terms to simplify the question. Then you bring all of the x to one side by either subtracting or adding it. Combine the rest if necessary. Solve for x.

OpenStudy (anonymous):

do you understand what you need to do ?